# Determination of the density in the linear elastic wave equation

### v2

@@ 4. S-waves and proof of the main theorem
$=$ $g^{jm}\Gamma^{\beta}_{jm}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\alpha}a_{0\beta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\beta}a_{0\alpha}+g^{jm}\Gamma^{3}_{jm}\mu\partial_{3}a_{0\alpha}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ ${g^{33}\Gamma^{\beta}_{33}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}}+{g^{\eta\varepsilon}\Gamma^{\beta}_{\eta\varepsilon}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}}+\mu(g^{33}\Gamma^{\beta}_{33}+g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})$
$+(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu\partial_{3}a_{0\alpha}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Collecting $\mathcal{O}(1)$ terms in equation (6) with $i=\alpha$, we obtain
(14) $\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)=\rho H,$
@@
$H=$ ${(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-{(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}}$
$+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$-2c_{P}^{2}c_{S}^{-1}\partial_{\alpha}c_{S}\frac{1}{c_{P}^{2}-c_{S}^{2}}\left((c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\right)$
$+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\left({(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\right)$
$+\partial_{\alpha}\log\sqrt{\rho}\frac{2(c_{P}^{2}-2c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\left({\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}}+g^{\gamma\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\gamma}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}-{\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}a_{0\delta}}\right)$
@@
$-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}-\partial^{2}_{33}a_{0\alpha}-\rho^{-1}g^{\beta\gamma}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}-c_{S}^{2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}-g^{\beta\gamma}\rho^{-1}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}-{\mu g^{\beta\gamma}\partial_{\alpha}\partial_{\gamma}a_{0\beta}}$
$+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}+2g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}+\partial_{3}\rho g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+2c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\rho\Gamma^{\beta}_{\alpha\gamma}a_{0\beta}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+c_{S}^{2}(-c_{S}^{-1}g^{\beta\gamma}\partial_{\gamma}c_{S})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}({\partial_{\alpha}a_{0\beta}}+\partial_{\beta}a_{0\alpha})$
@@
$H=$ $-\partial^{2}_{33}a_{0\alpha}-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}$
$-g^{\beta\gamma}\rho^{-1}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}-c_{S}^{2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}-\rho^{-1}g^{\beta\gamma}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\gamma}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-(c_{P}^{2}-2c_{S}^{2})\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}$
$+3g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+4c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\gamma}a_{0\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}-\rho^{-1}\partial_{\alpha}\lambda c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}a_{0\delta}$
$-c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}a_{0\alpha}+\mu g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}a_{0\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+4c_{S}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}$
@@
$H=$ $-\partial^{2}_{33}A_{S}c_{S}\eta_{\alpha}{-2\partial_{3}A_{S}\partial_{3}(c_{S}\eta_{\alpha}})-c_{S}^{-1}\partial_{3}c_{S}^{2}\partial_{3}A_{S}\eta_{\alpha}-2\partial_{3}\log\sqrt{\rho}\partial_{3}A_{S}c_{S}\eta_{\alpha}$
${-2\partial_{3}(c_{S}\log\sqrt{\rho}A_{S}\partial_{3}\eta_{\alpha}})+g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}A_{S}c_{S}\eta_{\beta}$
$-g^{\beta\gamma}\partial_{\beta}c_{S}^{2}\partial_{\gamma}A_{S}c_{S}\eta_{\alpha}-2g^{\beta\gamma}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}A_{S}\eta_{\alpha}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\gamma}(c_{S}\eta_{\alpha}})$
$-\mu g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}A_{S}c_{S}\eta_{\alpha}-{2\mu g^{\beta\gamma}\partial_{\beta}A_{S}\partial_{\gamma}(c_{S}\eta_{\alpha}})$
$-g^{\beta\gamma}\partial_{\gamma}c_{S}^{2}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\beta}})$
$+{2g^{\beta\gamma}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}A_{S}\eta_{\gamma}}+{2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\gamma}})+2g^{\beta\gamma}c_{S}^{3}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\partial_{\beta}c_{S}^{2}A_{S}c_{S}\eta_{\gamma}+4\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\gamma}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\gamma}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}c_{S}A_{S}\eta_{\gamma}$
$-\lambda\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}\partial_{\alpha}g^{\delta\varepsilon}+3g^{\delta\gamma}\mu\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}A_{S}c_{S}\eta_{\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}A_{S}c_{S}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\gamma}$

### v3

@@ Appendix A Proof of lemmas
$=$ $g^{jm}\Gamma^{\beta}_{jm}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\varepsilon}g^{\varepsilon\delta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\alpha}a_{0\beta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\beta}a_{0\alpha}+g^{jm}\Gamma^{3}_{jm}\mu\partial_{3}a_{0\alpha}$
[…]
$=$ ${g^{33}\Gamma^{\beta}_{33}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\varepsilon}g^{\varepsilon\delta}}+{g^{\eta\varepsilon}\Gamma^{\beta}_{\eta\varepsilon}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\varepsilon}g^{\varepsilon\delta}}+\mu(g^{33}\Gamma^{\beta}_{33}+g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})$
$+(g^{33}\Gamma^{3}_{33}+g^{\beta\varepsilon}\Gamma^{3}_{\beta\varepsilon})\mu\partial_{3}a_{0\alpha}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Using equation (10) with $i=\alpha$, we obtain
(19) $\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)=\rho H_{\alpha},$
@@
$H_{\alpha}=$ ${(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\alpha}\partial_{\beta}a_{0\varepsilon}}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}$
$+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\varepsilon}-{(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\Gamma^{\delta}_{\beta\varepsilon}\partial_{\alpha}a_{0\delta}}$
$+2g^{\beta\varepsilon}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$-2c_{P}^{2}c_{S}^{-1}\partial_{\alpha}c_{S}\frac{1}{c_{P}^{2}-c_{S}^{2}}\left((c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}\right)$
$+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\left({(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}\right)$
$+\partial_{\alpha}\log\sqrt{\rho}\frac{2(c_{P}^{2}-2c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\left({\frac{\lambda+\mu}{\rho}g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}}+g^{\varepsilon\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\varepsilon}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}a_{0\varepsilon}-{\frac{\lambda+\mu}{\rho}g^{\beta\varepsilon}\Gamma^{\delta}_{\beta\varepsilon}a_{0\delta}}\right)$
@@
$-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}-\partial^{2}_{33}a_{0\alpha}-\rho^{-1}g^{\beta\varepsilon}\partial_{\beta}\mu\partial_{\varepsilon}a_{0\alpha}-c_{S}^{2}g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}a_{0\alpha}-g^{\beta\varepsilon}\rho^{-1}\partial_{\varepsilon}\mu\partial_{\alpha}a_{0\beta}-{\mu g^{\beta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\beta}}$
$+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}+2g^{\delta\varepsilon}c_{S}^{2}\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}a_{0\beta}+\partial_{3}\rho g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{0\beta}+2c_{S}^{2}g^{\delta\varepsilon}\partial_{\delta}\rho\Gamma^{\beta}_{\alpha\varepsilon}a_{0\beta}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+g^{\varepsilon\delta}\Gamma^{\beta}_{\alpha\varepsilon}c_{S}^{2}\partial_{\delta}a_{0\beta}+g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+c_{S}^{2}(-c_{S}^{-1}g^{\beta\varepsilon}\partial_{\varepsilon}c_{S})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}({\partial_{\alpha}a_{0\beta}}+\partial_{\beta}a_{0\alpha})$
@@
$H_{\alpha}=$ $-\partial^{2}_{33}a_{0\alpha}-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}$
$-g^{\beta\varepsilon}\rho^{-1}\partial_{\beta}\mu\partial_{\varepsilon}a_{0\alpha}-c_{S}^{2}g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}a_{0\alpha}-\rho^{-1}g^{\beta\varepsilon}\partial_{\varepsilon}\mu\partial_{\alpha}a_{0\beta}$
$+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\varepsilon}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\varepsilon\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\varepsilon}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}-(c_{P}^{2}-2c_{S}^{2})\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}$
$+3g^{\delta\varepsilon}c_{S}^{2}\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}a_{0\beta}+g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{0\beta}+4c_{S}^{2}g^{\delta\varepsilon}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\varepsilon}a_{0\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}a_{0\varepsilon}-\rho^{-1}\partial_{\alpha}\lambda c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}a_{0\delta}$
$-c_{S}\partial_{\varepsilon}c_{S}g^{\beta\varepsilon}\partial_{\beta}a_{0\alpha}+\mu g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}a_{0\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+4c_{S}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\varepsilon}$
@@
$H_{\alpha}=$ $-\partial^{2}_{33}A_{S}c_{S}\eta_{\alpha}{-2\partial_{3}A_{S}\partial_{3}(c_{S}\eta_{\alpha}})-c_{S}^{-1}\partial_{3}c_{S}^{2}\partial_{3}A_{S}\eta_{\alpha}-2\partial_{3}\log\sqrt{\rho}\partial_{3}A_{S}c_{S}\eta_{\alpha}$
${-2\partial_{3}(c_{S}\log\sqrt{\rho}A_{S}\partial_{3}\eta_{\alpha}})+g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}\partial_{3}A_{S}c_{S}\eta_{\beta}$
$-g^{\beta\varepsilon}\partial_{\beta}c_{S}^{2}\partial_{\varepsilon}A_{S}c_{S}\eta_{\alpha}-2g^{\beta\varepsilon}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\varepsilon}A_{S}\eta_{\alpha}-{2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\varepsilon}(c_{S}\eta_{\alpha}})$
$-\mu g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}A_{S}c_{S}\eta_{\alpha}-{2\mu g^{\beta\varepsilon}\partial_{\beta}A_{S}\partial_{\varepsilon}(c_{S}\eta_{\alpha}})$
$-g^{\beta\varepsilon}\partial_{\varepsilon}c_{S}^{2}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}-{2g^{\beta\varepsilon}c_{S}^{2}\partial_{\varepsilon}\log\sqrt{\rho}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}}-{2g^{\beta\varepsilon}c_{S}^{2}\partial_{\varepsilon}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\beta}})$
$+{2g^{\beta\varepsilon}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}A_{S}\eta_{\varepsilon}}+{2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\varepsilon}})+2g^{\beta\varepsilon}c_{S}^{3}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$+2\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\varepsilon\beta}\partial_{\beta}c_{S}^{2}A_{S}c_{S}\eta_{\varepsilon}+4\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\varepsilon\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\varepsilon}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}A_{S}c_{S}\eta_{\varepsilon}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}A_{S}c_{S}\eta_{\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}c_{S}A_{S}\eta_{\varepsilon}$
$-\lambda\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}\partial_{\alpha}g^{\delta\varepsilon}+3g^{\delta\varepsilon}\mu\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}A_{S}c_{S}\eta_{\beta}+g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}\mu\partial_{\delta}A_{S}c_{S}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\varepsilon}$